distance through a point(origin)

if y=[6;-3;7] and u=[-6,-2,-3] compute the distance d , from the line through u and the origin. plz show work...d=?
Answer #1
equation of line
(x - 0)/(-6 - 0) = (y-0)/(-2-0) = (z-0)/(-3-0)

x/6 = y/2 = z/3 = k(say)

so (6k,2k,3k) is a general point

6-6k , -3-2k , 7-3k is vector
making dot product zero with line
k = 51/49

distance = 6.396 units

answered by: Janessa